OpenStudy (anonymous):

Let C represent the total cost (in dollars), and let S represent the amount of sugar (in tons) transported. Write an equation relating C to S . Then use this equation to find the total cost to transport 14 tons of sugar.

OpenStudy (anonymous):

The Sugar Sweet Company is going to transport its sugar to market. It will cost $3500 to rent trucks, and it will cost an additional $250

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) Here at OpenStudy, we like to lead the user to the answer. Do you have an idea of where to start, dear? OpenStudy Ambassador; Compassionate.

OpenStudy (anonymous):

No

OpenStudy (perl):

I dont see how we can relate cost to the number of sugar tons

OpenStudy (anonymous):

i need an equation and total cost to transport 14 tons of sugar

OpenStudy (compassionate):

Hi, """Let C represent the total cost (in dollars), and let S represent the amount of sugar (in tons) transported. Write an equation relating C to S . Then use this equation to find the total cost to transport 14 tons of sugar.""" Lets begin by dissecting your question. >S represents the amount of sugar >C is the total cost. >ind the total cost to transport 14 tons. Okay, so we need the price. Therefore, we're solving for C. _____________________________________________________________ S = C. But wait, this can't be correct. We're missing some more information. """"The Sugar Sweet Company is going to transport its sugar to market. It will cost $3500 to rent trucks, and it will cost an additional $250"""" >3,500 to rent trucks and an additional 250. We know how: S(3,500) + 250 = C. Alright, so we know S is in tons. So, we're asked to find: the total for 14 tons. ______________________________________________________________ Our next step is to plug in 14 to S. 14(3,500) + 25 = C Now, just simplify to solve for C. OpenStudy Ambassador; Compassionate.

OpenStudy (anonymous):

what is the equation

OpenStudy (anonymous):

i cant put that equation

OpenStudy (compassionate):

S(3,500) + 250 = C. The equation may be in a different form. Look for on like that.